# Smooth surface reconstruction of earthquake faults from distributed moment-potency-tensor solutionsC

### v2

@@ 4 Derivation of the optimal elevation field $\Gamma_{*}$
We have introduced $P(\Gamma|{\bf a})$ by recognizing the subjective nature of the mapping from two degrees of freedom in the expected normal $n_{*}$ to one degree of freedom in the smooth surface elevation. With the prior distribution $P(\Gamma|{\bf a})$ exemplified by eq. (14), we derive an analytical expression of the most probable surface estimate $\Gamma_{*}({\bf a})$ given the potency $\bf a$}. In the previous section, $P(\Gamma|{\bf a})$ was constructed exclusively based on the $n$-vector identity. This formulation is equivalent to a free boundary condition, where the partial differential equation is solved without external boundary constraints. We will mainly analyze this case. A generalization to cases with imposed boundary conditions will be addressed at the end of this section.
@@ 4.1 Governing equation
Hereafter, the local basis vectors associated with $(\theta,\phi)$ on the unit hemisphere aligned with $\bf n$} are denoted as $\bf n$} (radial), ${\bf e}_{\theta}$ (polar), and ${\bf e}_{\phi}$ (azimuthal).
@@ 4.2 Solving a stochastic partial differential equation for fault elevation
The geometry $\Gamma$ of a smooth surface is equivalent to both the elevation field $z$ and the normal-vector field $n$, but formally, eq. (7) can be read as a partial differential equation of $z$ with a random external force $(n_{\perp}/n_{z}){\bf n}_{\parallel}$ imposed. The probability distribution $P(\Gamma|{\bf a};\kappa)$ (eq. 14) describes the fluctuations of the $n$-vector around its expectation $n_{*}$. Here, we derive an analytical solution of this stochastic partial differential equation of eqs. (7) and (14). Strictly speaking, the presence of significant non-integrable components (eq. 12) of $n_{*}$ necessitates an explicit operation to remove them. However, since the results of this operation depend on the details of the prior $P(\Gamma|{\bf a};\kappa)$, we consider such scenarios out of scope for the present analysis to maintain generality. The full-order derivation addressing this complexity is provided in Appendix C.
Although eq. (14) is a central-limit distribution, it is non-Gaussian due to the $n$-vector normalization, not solvable via linear operations. The explicit form of this distribution is obtained by writing the $n$-vector as a function of $\nabla z$ (eq. 37 in the next section) in eq. (14). In the dislocation limit, those non-Gaussian fluctuations become infinitesimal, and the distribution reduces to a Gaussian peaked at $\Gamma_{*}$, which we evaluate below.
@@
Here, $\bar{z}$ denotes the expected elevation field, ${\bf w}(\cdot,\cdot)$ represents the inverse covariance, and $\mathcal{Z}^{\prime\prime}$ represents normalization. Note that boundary conditions exceeding the determination of the vertical translation introduce further overdeterminacy, whereby the solution varies according to the assigned weights of boundary penalties; this is because the solution of the governing equation (eq. 20), $z(x,y)=z(0,0)-\int_{(0,0)}^{(x,y)}d{\bf l}\cdot\frac{n_{\perp}}{n_{z}}{\bf n}_{\parallel}$, is unique up to the vertical offset $z(0,0)$. The quadratic form in eq. (36) includes forms of $P(\Gamma|{\bf a})$ other than eq. (14) when the higher orders are dropped in the dislocation limit. This generalized probability density function eq. (36) is applicable also to surface reconstruction from the point clouds of microseismicity focal mechanisms. In our teleseismic application, we use eq. (36) for computational acceleration (§6.2).
@@ 5 Surface reconstruction from synthetic noisy normal vectors
We treat two model shapes. The first shape is a uniaxial twist (§5.1). This shape is characterized by biaxial slopes, providing a simplest instance of a torn surface (i.e. eq. 11 violation) if a uniaxial interpolation is applied to the $n$-vector field. The uniaxial twist also represents a developable surface, the surface that can be exactly expressed by spline interpolation without error. These make this geometry a useful benchmark for comparison with the existing spline-based approach (Shimizu et al., 2021). The second is a bowl shape (§5.2), an example of non-developable surfaces that intrinsically require stochastic reconstruction for accurate reconstruction. This geometry resembes the earthquake fault estimated in our case study of the 2013 Balochistan earthquake (§6.3).
@@ 5.3 Robustness tests of surface reconstruction
Figure 8 presents a summary of the errors and residuals for the two tests described above across varying noise levels. The estimation errors $1-\hat{n}\cdot n_{0}$ and data residuals $1-\hat{n}\cdot n_{*}$ for normal vectors (Fig. 8a, b) both scale linearly with the noise amplitude $1-n_{0}\cdot n^{*}$. Notably, the data residuals are visibly smaller than the noise amplitude, indicating significant variance reduction. The estimation error $1-\hat{n}\cdot n_{0}$ is smaller than the data residual $1-\hat{n}\cdot n_{*}$. This suggests that the the reconstruction process does not merely explain the noisy input $n$-vectors but rather effectively denoises them.
@@ 7 Conclusion
We have investigated surface reconstruction from the density field of the potency tensor, where a smooth surface is constructed from the $n$-vector field indicated by the nodal planes of distributed potency-density-tensor solutions. The practical contribution of this work is the derivation of a formula that enables the three-dimensional surface reconstruction unaddressed by Shimizu et al. (2021). A key finding is that the problem of earthquake fault reconstruction is overdetermined, thus necessitating a probabilistic, Bayesian approach: a criterion is necessary to reduce the two degrees of freedom of the surface $n$-vector to the single degree of freedom of the surface elevation $z$. This overdetermination is a byproduct of the potency’s capacity to capture the inelastic earthquake source beyond displacement discontinuity. The $n\mapsto z$ mapping is a dimensionality reduction, entrusting the definition of what constitutes faulting to our subjectivity to discern the true nature of the earthquake fault. We have introduced the prior that treats an earthquake fault as a smooth interface embedded in a three-dimensional space, compatible with linear continuum mechanics, which will provide one normal form as in slip inversion. The same principle applies to surface reconstruction from distributed point sources, and our formulation is directly applicable to such cases.
@@
Because the degrees of freedom of the source model at each coordinate are six for inelastic strain and five for displacement discontinuity, the PDTI estimating the potency density is higher-dimensional than slip inversion estimating the displacement discontinuity. The underlying physics is eq. (2), where inelastic strain is represented as the tensor product of two vectors. Equation (2) signifies the compatibility condition for volume-less areal source distributions, serving as a constraint when projecting inelastic strain onto displacement discontinuity. While this constraint appears as a determinant-free condition for point sources, it is the regularity condition of the surface indicated by the $n$-vector field (eq. 11) for areal sources. The well-established fact in point-source analysis that a seismic source cannot be exhaustively expressed as a displacement discontinuity has been largely missed in finite-source analyses since the advent of slip inversion. However, this reappears in the PDTI as the overdetermined nature of surface reconstruction, where the areal source cannot be fully described by a displacement discontinuity. The optimal model in the PDTI is generally not a displacement discontinuity even after the earthquake fault is reconstructed. A research field of continuum theory originating with seismological descriptions of point sources, the inference of incompatible inelastic strains in linear continua is thus formalized as the PDTI for areal sources. It accesses a latent dimension of information that dislocation theory has left unaddressed by its nature, and our reconstruction of a smooth surface is but the first step toward decoding the apparent incompatibility inherent in the areal source representation.
@@
Having elucidating the above essence of surface reconstruction, we have formulated and solved the problem of stochastic surface reconstruction from a given noisy $n$-vector field. Our synthetic tests verify that the derived solution successfully reproduces the original surface from noisy $n$-vector data. Even for narrow faults, our three-dimensional probabilistic approach can better reproduce the original shape than previously proposed deterministic quasi-two-dimensional approaches. A preliminary analysis of the 2013 Balochistan earthquake captures along-dip shape variability that was not resolved in Shimizu et al. (2021), which validates our proposal. Neither the prior to reconstruct a fault as a smooth surface nor our recipe to perturbatively reconstruct a surface from the initial guess is yet exhaustive. Nonetheless, they establish a reference framework for earthquake fault reconstruction from the measured seismic moment.
@@ Acknowledgements.
Waveform data inverted in our application were downloaded from the Seismological Facility for the Advancement of Geoscience Data Management Center, processed in Shimizu et al. (2020). The ground topography in this paper refers to SRTMGL1 (NASA JPL, 2013).
Appendix A Derivation of Eq. (2) as the compatibility condition for areal inelastic strain fields
Equation (2) represents an integrability condition for inelastic strain in the limit of infinitesimal volume, which does not necessarily hold for volumetric sources even when thickness is ignored. Since the role of eq. (2) as the integrability condition, the compatibility condition in continuum mechanics, is seldom explicitly articulated in the literature, we provide its derivation here to clarify the distinction between inelastic strain and the displacement-discontinuity representation.
The complementarity condition for a symmetric inelastic strain $\epsilon^{*}$ to possess a corresponding displacement field is given by (Saada, 2013)
@@
where $e_{ikr}$ is the permutation symbol, and $\partial$ denotes partial differentiation. While the compatibility condition is typically applied to elastic strain, it remains valid for inelastic deformation as long as the inelastic strain is defined as the symmetrized gradient of an inelastic displacement field.
We apply this compatibility condition to a distribution of inelastic strain with thickness $\Delta h$ along a pseudo boundary $\Gamma$. For $\boldsymbol{\xi}\in\Gamma$, we consider an asymptotic representation for a small $\Delta h$ while keeping the areal potency density $D:=\int^{\Delta h/2}_{-\Delta h/2}dh\epsilon^{*}(\boldsymbol{\xi}+h\mathbf{n}(\boldsymbol{\xi}))$ constant. In this asymptotics, the gradient in the $\bf n$}-direction scales as $\Delta h^{-1}$. Consequently, the triple integral of eq. (A.2) along the $\bf n$}-direction reduces to:
@@
Finally, by evaluating the symmetric double cross product with $\bf n$}, ${\bf n}\times l.h.s.\times{\bf n}=\mathcal{O}(\Delta h)$, we obtain the following rewritten form of the compatibility (Sato and Moore, 2022):
@@
Note that eq. (2) merely describes the incompatibility with displacement discontinuity localized on a pseudo-boundary $\Gamma$, which is a surface that needs to be smooth to maintain consistency with the continuum mechanics. In the context of surface reconstruction, this manifests as eq. (11). Crucially, the derivation above does not exclude the possibility that a physically compatible inelastic strain field may appear seemingly incompatible when mapped onto a single smooth surface $\Gamma$. A violation of eq. (2) represents the potential departure of the actual source geometry from the model space expressed by the limit of a single smooth surface $\Gamma$.
@@ Appendix C Correction of $\Gamma_{*}$ in evaluating non-integrable potency components
$+\left[\sin\theta_{*}^{\prime}(\phi-\phi*^{\prime})+\cos\theta_{*}^{\prime}(\theta-\theta_{*}^{\prime})(\phi-\phi_{*}^{\prime})\right]{\bf e}_{\phi}^{*\prime}.$
@@
$-\left[\sin\theta_{*}^{\prime}(\phi-\phi*^{\prime})+\cos\theta_{*}^{\prime}(\theta-\theta_{*}^{\prime})(\phi-\phi_{*}^{\prime})\right]({\bf n}_{*}\cdot{\bf e}_{\phi}^{*\prime}).$
@@
$-2[\sin\theta_{*}^{\prime}(\phi-\phi*^{\prime})+\cos\theta_{*}^{\prime}(\theta-\theta_{*}^{\prime})(\phi-\phi_{*}^{\prime})]({\bf n}_{*}\cdot{\bf e}_{\phi}^{*\prime})\}.$

### v3

@@ 4 Derivation of the optimal elevation field $\Gamma_{*}$
We have introduced $P(\Gamma|{\bf a})$ by recognizing the subjective nature of the mapping from two degrees of freedom in the expected normal $n_{*}$ to one degree of freedom in the smooth surface elevation. With the prior distribution $P(\Gamma|{\bf a})$ exemplified by eq. (14), we derive an analytical expression of the most probable surface estimate $\Gamma_{*}({\bf a})$ given the potency $\bf a$}. In the previous section, $P(\Gamma|{\bf a})$ was constructed exclusively based on the $n$-vector identity. This formulation is equivalent to a free boundary condition, where the partial differential equation is solved without external boundary constraints. We mainly analyze this case. A generalization to cases with imposed boundary conditions will be addressed at the end of this section.
@@ 4.1 Governing equation
Hereafter, the local basis vectors associated with $(\theta,\phi)$ on the unit hemisphere associated with $\bf n$} are denoted as $\bf n$} (radial), ${\bf e}_{\theta}$ (polar), and ${\bf e}_{\phi}$ (azimuthal).
@@ 4.2 Solving a stochastic partial differential equation for fault elevation
The geometry $\Gamma$ of a smooth surface is equivalent to both the elevation field $z$ and the normal-vector field $n$, but formally, eq. (7) can be read as a partial differential equation of $z$ with a random external force $(n_{\perp}/n_{z}){\bf n}_{\parallel}$ imposed. The probability distribution $P(\Gamma|{\bf a};\kappa)$ (eq. 14) describes the fluctuations of the $n$-vector around its expectation $n_{*}$. Here, we derive an analytical solution of this stochastic partial differential equation of eqs. (7) and (14). Strictly speaking, the presence of significant non-integrable components (eq. 12) of $n_{*}$ necessitates an explicit operation to remove them. However, since the results of this operation depend on the details of the prior $P(\Gamma|{\bf a};\kappa)$, we consider such scenarios out of scope for the present analysis to maintain generality. We focus on the case where $n\simeq n_{*}$ holds such that the functional form of the prior is irrelevant. The full-order derivation addressing the higher-order complexity is provided in Appendix C.
Although eq. (14) is a central-limit distribution, it is non-Gaussian due to the $n$-vector normalization, not solvable via linear operations. The explicit form of this distribution is obtained by writing the $n$-vector as a function of $\nabla z$ (eq. 37 in the next section) in eq. (14). However, in the dislocation limit, those non-Gaussian fluctuations become infinitesimal, and the distribution reduces to a Gaussian peaked at $\Gamma_{*}$, which we evaluate below.
@@
Here, $\bar{z}$ denotes the expected elevation field, ${\bf w}(\cdot,\cdot)$ is the inverse covariance, and $\mathcal{Z}^{\prime\prime}$ represents normalization. Note that boundary conditions exceeding the determination of the vertical translation introduce further overdeterminacy, whereby the solution varies according to the assigned weights of boundary penalties; this is because the solution of the governing equation (eq. 20), $z(x,y)=z(0,0)-\int_{(0,0)}^{(x,y)}d{\bf l}\cdot\frac{n_{\perp}}{n_{z}}{\bf n}_{\parallel}$, is unique up to the vertical offset $z(0,0)$. The quadratic form in eq. (36) includes forms of $P(\Gamma|{\bf a})$ other than eq. (14) when the higher orders are dropped in the dislocation limit. This generalized probability density function eq. (36) is applicable also to surface reconstruction from the point clouds of microseismicity focal mechanisms. In our teleseismic application, we use eq. (36) for computational acceleration (§6.2).
@@ 5 Surface reconstruction from synthetic noisy normal vectors
We treat two model shapes. The first shape is a uniaxial twist (§5.1). This shape is characterized by biaxial slopes, providing a simplest instance of a torn surface (i.e. eq. 11 violation) if a uniaxial interpolation is applied to the $n$-vector field. The uniaxial twist also represents a developable surface, the surface that can be exactly expressed by spline interpolation without error. These make this geometry a useful benchmark for comparison with the existing spline-based approach (Shimizu et al., 2021). The second is a bowl shape (§5.2), an example of non-developable surfaces that intrinsically require stochastic reconstruction for accurate reconstruction. This geometry resembles the earthquake fault estimated in our case study of the 2013 Balochistan earthquake (§6.3).
@@ 5.3 Robustness tests of surface reconstruction
Figure 8 presents a summary of the errors and residuals for the two tests described above across varying noise levels. The estimation errors $1-\hat{n}\cdot n_{0}$ and data residuals $1-\hat{n}\cdot n_{*}$ for normal vectors (Fig. 8a, b) both scale linearly with the noise amplitude $1-n_{0}\cdot n^{*}$. Notably, the data residuals are visibly smaller than the noise amplitude, indicating significant variance reduction. The estimation error $1-\hat{n}\cdot n_{0}$ is smaller than the data residual $1-\hat{n}\cdot n_{*}$. This suggests that the reconstruction process does not merely explain the noisy input $n$-vectors but rather effectively denoises them.
@@ 7 Conclusion
We have investigated surface reconstruction from the density field of the potency tensor, where a smooth surface is constructed from the $n$-vector field indicated by the nodal planes of distributed potency-density-tensor solutions. The practical contribution of this work is the derivation of a formula that enables the three-dimensional surface reconstruction unaddressed by Shimizu et al. (2021). A key finding is that the problem of earthquake fault reconstruction is overdetermined, thus necessitating a probabilistic, Bayesian approach: a criterion is necessary to reduce the two degrees of freedom of the surface $n$-vector to the single degree of freedom of the surface elevation $z$. This overdetermination is a byproduct of the potency’s capacity to capture the inelastic earthquake source beyond displacement discontinuity. The $n\mapsto z$ mapping is a dimensionality reduction, entrusting the definition of what constitutes faulting to our subjectivity to discern the true nature of the earthquake fault. We have introduced the prior that treats an earthquake fault as a smooth interface embedded in a three-dimensional space, compatible with a dislocation within a linear continuum, which will provide one normal form as in slip inversion. The same principle applies to surface reconstruction from distributed point sources, and our formulation is directly applicable to such cases.
@@
Because the degrees of freedom of the source model at each coordinate are six for inelastic strain and five for displacement discontinuity at most, the PDTI estimating the potency density is higher-dimensional than slip inversion estimating the displacement discontinuity. The underlying physics is eq. (2), where inelastic strain is represented as the tensor product of two vectors. Equation (2) signifies the requirement for volume-less areal source distributions, serving as a constraint when projecting inelastic strain onto displacement discontinuity. While this constraint is a simple determinant-free condition for point sources, it raises the regularity condition of the surface indicated by the $n$-vector field (eq. 11) for areal sources. The well-established fact in point-source analysis that a seismic source cannot be exhaustively expressed as a displacement discontinuity has been largely missed in finite-source analyses since the advent of slip inversion. However, this reappears in the PDTI as the overdetermined nature of surface reconstruction, where the areal source cannot be fully described by a displacement discontinuity. The optimal model in the PDTI is generally not a displacement discontinuity even after the earthquake fault is reconstructed. A research field of continuum theory originating with seismological descriptions of point sources, the inference of inelastic strains with excess degrees of freedom is thus formalized as the PDTI for areal sources. Our reconstruction of a smooth surface is but the first step toward decoding the volumetric rupture episodes compressed into the excess degrees of freedom in the areal source representation.
@@
Having elucidated the above essence of surface reconstruction, we have formulated and solved the problem of stochastic surface reconstruction from a given noisy $n$-vector field. Our synthetic tests verify that the derived solution successfully reproduces the original surface from noisy $n$-vector data. Even for narrow faults, our three-dimensional probabilistic approach can better reproduce the original shape than previously proposed deterministic quasi-two-dimensional approaches. A preliminary analysis of the 2013 Balochistan earthquake captures along-dip shape variability that was not resolved in Shimizu et al. (2021), which validates our proposal. Neither the prior to reconstruct a fault as a smooth surface nor our recipe to perturbatively reconstruct a surface from the initial guess is yet exhaustive. Nonetheless, they establish a reference framework for earthquake fault reconstruction from the measured seismic moment.
@@ Acknowledgements.
Waveform data inverted in our application were downloaded from the Seismological Facility for the Advancement of Geoscience Data Management Center and were processed in Shimizu et al. (2020). The ground topography in this paper refers to SRTMGL1 (NASA JPL, 2013).
Appendix A Derivation of Eq. (2) from the compatibility condition for areal inelastic strain fields
Equation (2) represents an integrability condition for inelastic strain in the limit of infinitesimal volume, which does not necessarily hold for volumetric sources even when thickness is ignored. Since the role of eq. (2) as the integrability condition, often played by the compatibility condition in continuum mechanics, is seldom explicitly articulated in the literature, we provide its derivation here to clarify the distinction between inelastic strain and the displacement-discontinuity representation.
The compatibility condition for a symmetric inelastic strain $\epsilon^{*}$ to possess a corresponding displacement field is given by (Saada, 2013)
@@
where $e_{ikr}$ is the permutation symbol, and $\partial$ denotes partial differentiation.
While the compatibility condition is typically imposed on the elastic strain, it violates the compatibility condition in dislocation problems. An inelastic strain must be incompatible to act as a stress source; otherwise, it could simply be compensated for by a compatible elastic deformation without generating internal stress. To preserve the compatibility of the total strain, the elastic strain absorbs this incompatibility precisely where the inelastic source exists. For an inelastic source localized as a Dirac delta function, the violation of the compatibility equation is proportional to the second spatial derivative of the delta function. Here, the overall compatibility of elastic strains is recovered when integrated three times over space. By the same token, the areal inelastic strain must also satisfy the compatibility condition in the sense of a triple spatial integral.
We apply this weakened compatibility condition to a distribution of inelastic strain with thickness $\Delta h$ along a pseudo boundary $\Gamma$. For $\boldsymbol{\xi}\in\Gamma$, we consider an asymptotic representation for a small $\Delta h$ while keeping the areal potency density $D:=\int^{\Delta h/2}_{-\Delta h/2}dh\epsilon^{*}(\boldsymbol{\xi}+h\mathbf{n}(\boldsymbol{\xi}))$ constant. In this asymptotics, the gradient in the $\bf n$}-direction scales as $\Delta h^{-1}$. Consequently, the triple integral of eq. (A.2) along the $\bf n$}-direction reduces to:
@@
Finally, by evaluating the symmetric double cross product with $\bf n$}, ${\bf n}\times l.h.s.\times{\bf n}=\mathcal{O}(\Delta h)$, we obtain the following weakened form of the compatibility (Sato and Moore, 2022):
@@
Equation (2) describes the requirement for the limit of a displacement discontinuity localized on a pseudo-boundary $\Gamma$. To maintain consistency with continuum mechanics, the pseudo-boundary $\Gamma$ must be a smooth surface in this limit, a geometric constraint that manifests as eq. (11) in surface reconstruction. We remark that those requirements do not exclude the possibility that zero-volume inelastic strain fields, such as fault networks, may violate eqs. (2) and (11) when mapped onto a single smooth surface. A violation of these conditions implies the potential departure of the actual source geometry from the given model space.
@@ Appendix C Correction of $\Gamma_{*}$ in evaluating non-integrable potency components
$+\left[\sin\theta_{*}^{\prime}(\phi-\phi_{*}^{\prime})+\cos\theta_{*}^{\prime}(\theta-\theta_{*}^{\prime})(\phi-\phi_{*}^{\prime})\right]{\bf e}_{\phi}^{*\prime}.$
@@
$-\left[\sin\theta_{*}^{\prime}(\phi-\phi_{*}^{\prime})+\cos\theta_{*}^{\prime}(\theta-\theta_{*}^{\prime})(\phi-\phi_{*}^{\prime})\right]({\bf n}_{*}\cdot{\bf e}_{\phi}^{*\prime}).$
@@
$-2[\sin\theta_{*}^{\prime}(\phi-\phi_{*}^{\prime})+\cos\theta_{*}^{\prime}(\theta-\theta_{*}^{\prime})(\phi-\phi_{*}^{\prime})]({\bf n}_{*}\cdot{\bf e}_{\phi}^{*\prime})\}.$